The number of years for $1800 to grow to $9400 at 8% quarterly?
A=P(1+r/m)^tm
9400 = 1800 (1 + 0.08/4)^4t
5.22 = (1 = 0.08/4) ^4t
I don't know how get the value of t from the equation
Could you please help me.

Answer by stanbon(75887)

The number of years for $1800 to grow to $9400 at 8% quarterly?
A=P(1+r/m)^tm
9400 = 1800 (1 + 0.08/4)^4t
5.22 = (1 + 0.08/4) ^4t
5.22=(1.02)^(4t)
Take the log of both sides to get:
log5.22 =(4t)(log1.02]
4t=[log5.22]/[log1.02]
4t=83.4483983...
t=20.86 yrs.
Cheers,
Stan H.